### v1

# iConFormer: Dynamic Parameter-Efficient Tuning with Input-Conditioned Adaptation

Abstract
Transfer learning based on full fine-tuning (FFT) of the pre-trained encoder and task-specific decoder becomes increasingly complex as deep models grow exponentially. Parameter efficient fine-tuning (PEFT) approaches using adapters consisting of small learnable layers have emerged as an alternative to FFT, achieving comparable performance while maintaining high training efficiency. However, the inflexibility of the adapter with respect to input instances limits its capability of learning task-specific information in diverse downstream tasks. In this paper, we propose a novel PEFT approach, input-Conditioned transFormer, termed iConFormer, that leverages a dynamic adapter conditioned on the input instances. To secure flexible learning ability on input instances in various downstream tasks, we introduce an input-Conditioned Network (iCoN) in the dynamic adapter that enables instance-level feature transformation. To be specific, iCoN generates channel-wise convolutional kernels for each feature and transform it using adaptive convolution process to effectively capture task-specific and fine-grained details tailor to downstream tasks. Experimental results demonstrate that by tuning just 1.6% to 2.8% of the Transformer backbone parameters, iConFormer achieves performance comparable to FFT in monocular depth estimation and semantic segmentation, while outperforming it in image classification and instance segmentation. Also, the proposed method consistently outperforms recent PEFT methods for all the tasks mentioned above.
Figure 1: Quantitative comparison with full fine-tuning (FFT) and PEFT approaches. The top graph compares depth prediction errors on NYU-v2 (Silberman et al. 2012), while the bottom graph presents performance for semantic segmentation on ADE20K (Zhou et al. 2017) and instance segmentation on COCO (Lin et al. 2014). iConFormer consistently outperforms recent PEFT methods in all dense tasks and surpasses FFT in instance segmentation.
Introduction
As deep neural networks (DNNs) grow increasingly complex, transfer learning—fine-tuning pre-trained models with task-specific data for downstream tasks—has become a widely adopted solution across diverse applications, including image classification, semantic segmentation, and object detection, to name a few. For instance, the model consisting of the pre-trained encoder (Xie et al. 2022b; He et al. 2022a) and task-specific decoder is fine-tuned, achieving remarkable performance gain when compared to the model trained from scratch. However, training large, complex models separately for each task and dataset is highly inefficient. Recently, parameter efficient fine-tuning (PEFT) approaches (Zaken, Ravfogel, and Goldberg 2021; Qi et al. 2022; Hu et al. 2021) that maximize the efficiency in terms of training parameters have emerged as an alternative to the above-mentioned full fine tuning (FFT) methodologies, achieving competitive performance even with limited computing resources while simplifying the training processes and deployment.
This remarkable progress in vision tasks is primarily driven by approaches including prompt tuning (Lester, Al-Rfou, and Constant 2021) and adapter (Houlsby et al. 2019), which have been successfully applied in natural language processing (NLP) tasks. Visual prompt tuning (Jia et al. 2022) is the first study to explore the potential of prompt tuning in visual recognition tasks, laying the foundation for the prompt tuning in the field of computer vision. In addition, the adapter-based PEFT methods (Chen et al. 2022; Yin et al. 2023) achieve significant training efficiency by applying the adapter used in NLP to the Vision Transformer (ViT) and its variants (Dosovitskiy et al. 2021; Liu et al. 2021, 2022).
While most PEFT-based approaches yield comparable performance to baseline methods using the FFT in the image classification task, they do not yet provide sufficiently satisfactory performance to compete with the FFT in other complex downstream tasks. Additionally, the scalability of prompt-based methods (Jia et al. 2022; Bahng et al. 2022) is significantly limited, leading to considerable performance degradation as the number of learnable parameters increases, as reported in (Chen et al. 2022). Adapter-based approaches (Chen et al. 2022; Hu et al. 2021; Yin et al. 2023) incorporate lightweight modules to reduce the number of trainable parameters while maintaining stable performance across a range of trainable parameter scales. However, the adapters always apply the same transformation to input features, ignoring individual characteristics of input instances. This may not be an issue when fully tuning DNNs, but it could be a limiting factor in improving performance in adapter-based PEFT methods. Namely, the inflexibility with respect to input instances exacerbates the transfer capability of adapter-based models with small learnable parts to downstream tasks, limiting their ability to capture unique and task-specific information.
Additionally, the Vision Transformer (ViT) backbone (Dosovitskiy et al. 2021) used in adapter-based models tends to focus on global information rather than fine-grained local details within an image. While this limitation can be partially addressed by employing the Swin Transformer (Liu et al. 2021), which utilizes local attention mechanisms, the constraint on the number of learnable parameters in PEFT still restricts the Swin Transformer’s capability to effectively capture local features. Consequently, this negatively affects the performance in dense prediction tasks that require fine-grained information.
To address the aforementioned issues, we propose a novel PEFT approach, input-Conditioned transFormer, termed iConFormer, that leverages a dynamic adapter where parameters are adjusted input instance-level. Unlike recent adapter-based approaches (Chen et al. 2022; Hu et al. 2021; Yin et al. 2023), iConFormer introduces an input-Conditioned Network (iCoN) that dynamically generates the parameters for each input feature in the adapter. This approach enables for effectively capturing task-specific and local details while keeping the number of learnable parameters small. The effectiveness of our method is evidenced by the quantitative analysis in Figure 1, which shows that iConFormer captures task-specific information more effectively than conventional methods. We also analyze the capability to capture fine-grained details by visualizing attention maps in Figure 3.
iConFormer achieves performance competitive to full fine-tuning (FFT) on both classification and dense prediction tasks including monocular depth estimation, semantic segmentation, and instance segmentation with only additional 1.6% to 2.8% backbone parameters. Our method even surpasses FFT for the image classification on CIFAR100 (Krizhevsky, Hinton et al. 2009) and the challenging instance segmentation task on COCO (Lin et al. 2014). Additionally, iConFormer also outperforms conventional PEFT methods for monocular depth estimation task on NYU-v2 (Silberman et al. 2012), demonstrating the effective utilization of pre-trained backbone parameters with additional learnable parameters dynamically finetuned for specific tasks. In summary, our contributions are threefold:
  • We propose iConFormer to enhance representation learning by dynamically adjusts only a small subset of parameters conditioned on input instances in the PEFT framework.
  • We demonstrate that iConFormer effectively captures fine-grained details with input-Conditioned Network (iCoN), leading to substantial improvements in dense prediction tasks.
  • Through comprehensive experiments on classification, monocular depth estimation, instance segmentation, and semantic segmentation, we show that iConFormer achieves remarkable performance by tuning only 1.6% to 2.8% of the Transformer backbone parameters.
Figure 2: Comparison of Full Fine-Tuning (FFT) and the proposed Parameter Efficient Fine-Tuning (PEFT) using iConFormer. (a) FFT, where all parameters are updated during training. (b) Our PEFT (iConFormer), where adapters sequentially are attached after the MLP layer in the Transformer. Inside the adapters, an Input-Conditioned Network (iCoN) generates input-conditioned convolutional kernels in a channel basis. By convolving features with these kernels, iCoN adaptively refines the features in accordance with the specific properties of the input, thereby enhancing the model’s capability to effectively process diverse data in the downstream tasks.
Related Work
Transformer in Vision.
Transformers, initially designed for Natural Language Processing (NLP) tasks such as machine translation (Vaswani 2017) and text generation (Devlin 2018), have achieved significant success in these areas. This success has led to a shift towards computer vision, starting with the Vision Transformer (ViT) (Dosovitskiy et al. 2021). Subsequently, various Transformer-based models (Liu et al. 2021; Xie et al. 2021; Carion et al. 2020; Lee et al. 2022, 2023; Ranftl, Bochkovskiy, and Koltun 2021; Hong et al. 2022) have achieved notable advancements in tasks including image classification (Krizhevsky, Sutskever, and Hinton 2012), semantic segmentation (Long, Shelhamer, and Darrell 2015; Chen et al. 2017; Cho et al. 2024; Kim, Choi, and Min 2022), object detection (Girshick et al. 2014; Ren et al. 2016; Redmon et al. 2016), image restoration (Dong et al. 2015; Liu, Breuel, and Kautz 2017), and depth estimation (Godard et al. 2019; Choi et al. 2021). Furthermore, transformers have significantly advanced vision recognition through large-scale pretraining (Chen, Xie, and He 2021; Caron et al. 2021; He et al. 2022a). However, their larger size compared to previously prevalent CNN backbones presents challenges for fine-tuning on specific tasks. In this context, our work explores methods to adapt pre-trained transformers into target tasks in a more effective and efficient way.
Parameter Efficient Fine Tuning.
Parameter Efficient Fine-Tuning (PEFT) methods enable the adaptation of large self-supervised pre-trained models (Xie et al. 2022c; He et al. 2022b; Yi et al. 2022; Choi et al. 2024a, b) to specific tasks without the need to train the entire model. In Natural Language Processing (NLP), notable approaches include adapter methods (Houlsby et al. 2019), which integrate small modules into the model while keeping the pre-trained parameters frozen, with only the added modules being fine-tuned. Additionally, other methods involve tuning specific components, such as bias or normalization layers (Zaken, Ravfogel, and Goldberg 2021; Qi et al. 2022), utilizing learnable prompt tokens (Lester, Al-Rfou, and Constant 2021), or applying low-rank matrix approximations (Hu et al. 2021) to update parameters more efficiently.
In computer vision, PEFT techniques inspired by NLP have shown significant progress. For instance, VPT (Jia et al. 2022) is the first method to apply prompt tuning approaches to visual classification tasks. AdaptFormer (Chen et al. 2022) employs a parallel adapter framework to enhance the effectiveness of parameter efficient fine-tuning for visual classification. KAdaptation (He et al. 2023) optimizes the adapter using Kronecker products. In addition, LoRand (Yin et al. 2023) employs multi-branch low-rank adapters to achieve impressive performance on dense prediction tasks. Our work focuses on adapter methods, proposing new approaches that aim to achieve performance comparable to or exceeding that of full fine-tuning across various downstream tasks.
Preliminary
Vision Transformer and its Variants
Vision Transformer (ViT) (Dosovitskiy et al. 2021), modified from the Transformer (Vaswani 2017) proposed in NLP, integrates image patches and positional encodings to capture spatial information. It consists of a patch embedding layer and multiple sequential encoder blocks, as depicted in Fig. 2(a). Given a batch of images $x\in\mathbb{R}^{B\times H\times W\times 3}$, the patch embedding layer transforms $x$ into sequential patches $x_{p}\in\mathbb{R}^{B\times M\times(P^{2}C)}$, where $H\times W$ is an image resolution, and $B$ is a batch size. $P\times P$ is the resolution of an image patch, $C$ is the output channel, and $M=HW/P^{2}$ is the number of image patches. The patches are linearly embedded into $D$ dimensions to generate the final input $x_{in}\in\mathbb{R}^{B\times M\times D}$.
In the Transformer encoder block, $x_{in}$ is first normalized using LayerNorm (Ba, Kiros, and Hinton 2016), and then processed by a multi-head self-attention layer (MHSA). The output is combined with $x_{in}$ via a residual connection:
$\acute{x_{in}}=\textrm{Attention}(\textrm{LN}(x_{in}))+x_{in}\,.$ (1)
Next, $\acute{x_{in}}$ is normalized and passed through the MLP layer, followed by residual connection:
$\tilde{x}=\textrm{MLP}(\textrm{LN}(\acute{x_{in}}))\,,$
$x_{out}=\tilde{x}+\acute{x_{in}}\,.$ (2)
This process is repeated $N$ times in the encoder block. In ViT, the self-attention mechanism captures global features by evaluating relationships between all image patches, enabling a comprehensive understanding of complex dependencies. Advancing this approach, Swin Transformer (Liu et al. 2021, 2022) introduces hierarchical attention with shifted windows, which enhances both computational efficiency and feature representation. Other variants (Wang et al. 2021; Xie et al. 2021; Li et al. 2023; Xia et al. 2024) leverage multi-scale feature extraction for specific vision tasks, demonstrating the adaptability of Transformer models to diverse computer vision challenges.
PEFT Methods
Parameter efficient fine-tuning (PEFT) methods, such as prompt tuning (Lester, Al-Rfou, and Constant 2021; Jia et al. 2022), low-rank adaptation (Hu et al. 2021), and adapters (Houlsby et al. 2019; Chen et al. 2022), are designed to reduce the number of trainable parameters needed for fine-tuning large models. Here, we briefly review adapter-based PEFT methods related to our approach, which will be detailed in the following section. Adapters introduce small, trainable modules between the layers of the pre-trained model. For instance, if the original model’s transformation is defined as $h=W\cdot x$, adapters modify this operation to $h=W\cdot x+A\cdot g(B\cdot x)$, where $B$ represents the down-projection of the input features, $A$ indicates the up-projection back to the original space, and $g(\cdot)$ denotes a non-linearity. These approaches allow for effective task adaptation while minimizing the number of parameters that need to be updated.
Figure 3: Comparison of Attention Maps from AdaptFormer and iConFormer. We visualize the attention maps using attention rollout (Abnar and Zuidema 2020). The top row represents input images, and the middle and bottom rows present the attention maps generated by AdaptFormer (Chen et al. 2022) and iConFormer, respectively, with both using Swin Transformer backbone (Liu et al. 2021). iConFormer more accurately delineates object regions and captures fine-grained semantics, compared to the AdaptFormer.
Proposed Method
Conventional adapter-based methods (Houlsby et al. 2019; Chen et al. 2022) rely on static parameter tunings, which may constrain their ability to capture diverse input features with high precision. To address these limitations, we propose iConFormer, a novel framework incorporating the input-Conditioned Network (iCoN), as illustrated in Figure 2 (b). Unlike conventional adapters, iConFormer dynamically generates convolutional kernels tailored to the unique characteristics of each input. This dynamic generation facilitates more accurate and flexible feature extraction, introducing locality inductive biases into the pre-trained Transformer encoder. This approach enables the model to effectively capture both global and local features within a parameter-efficient framework. Consequently, iConFormer not only improves the model’s performance but also ensures a more precise adaptation to varying input conditions.
Input-Conditioned Network (iCoN)
The input-Conditioned Network (iCoN) is a key component of the iConFormer framework. Inspired by the concept of dynamic filter networks (Jia et al. 2016), iCoN employs a channel-wise mechanism to generate convolutional kernels that are tailored to the unique characteristics of each input. This approach enhances parameter efficiency while maintaining the flexibility needed to capture diverse features effectively. Formally, the iCoN module generates channel-wise convolutional kernels using input features from the MLP output, denoted as $\tilde{x}\in\mathbb{R}^{B\times M\times D}$ of (2). First, the feature $\tilde{x}$ is down-projected to $d$ channels and reshaped from $M$ into the spatial dimensions $\hat{H}\times\hat{W}$, which correspond to $H/P\times W/P$. This reshaping effectively rearranges the patches into a spatial grid to produce the feature $\hat{x}\in\mathbb{R}^{B\times\hat{H}\times\hat{W}\times d}$. Subsequently, the iCoN module generates dynamic convolutional kernels from the reshaped feature $\hat{x}$. This process is mathematically represented as:
$\hat{x}=\text{Reshape}(\text{Down}(\tilde{x}))\,,$ (3)
$W_{dynamic}=\text{iCoN}(\hat{x})\,,$ (4)
where $W_{dynamic}\in\mathbb{R}^{B\times D\times K\times K}$ is the dynamically generated kernel and $K$ is the kernel size. For our implementation, we set $K$ to 3 and $d$ to 64 ($d\ll D$). The dynamically generated convolution kernel $W_{dynamic}$ is then applied to feature $\hat{x}$, enabling a more precise capture of each input’s distinct features. Afterward, the feature is reshaped back to $\mathbb{R}^{B\times M\times d}$, followed by applying a non-linear activation function and up-projection to restore the channel dimension to the original size $D$. This process produces the final output of adapter as follows:
$x_{A}=\text{Up}(\sigma(\text{Reshape}(\hat{x}\otimes W_{dynamic})))\,,$ (5)
[…]
$x_{out}=\gamma\cdot x_{A}+\acute{x_{in}}\,,$ (6)
where $\gamma$ is a weight that adjusts the impact of the adapter features. This weight is a learnable scalar, optimized during training.
By leveraging dynamically generated convolutional kernels, iConFormer refines the capture of input-specific features compared to conventional adapters, while maintaining parameter efficiency. This efficiency is achieved by employing a consistent set of parameters within the iCoN to generate convolutional kernels, even though the kernels are dynamically adapted for each specific input. Furthermore, the efficiency is enhanced by generating these dynamic kernels on a channel-wise basis, thereby minimizing the number of tunable parameters while preserving effective adaptation.
Consequently, iConFormer strikes a balance between effective feature extraction and computational efficiency relative to conventional PEFT methods, addressing their limitations and providing a robust solution for parameter-efficient adaptation.
Visual Analysis of Local Representation
To evaluate the performance of capturing both local and global information, Fig. 3 visualizes the attention maps that offer insights on where the model focuses on. The attention maps are generated using attention rollout (Abnar and Zuidema 2020) with AdaptFormer (Chen et al. 2022) and iConFormer, employing Swin Transformer (Liu et al. 2021) as the backbone. Attention rollout computes token attentions by recursively multiplying attention matrices across layers, revealing how attention is distributed across different regions of the image. AdaptFormer adopts a standard pipeline of conducting down-projection, non-linear activation, and up-projection using static weight parameters that are not dynamically adjusted conditioned on input features. While the attention maps of AdaptFormer exhibit ambiguous and scattered attention distributions, with limitations in capturing fine-grained local features such as object edges, the attention maps of iConFormer are significantly more focused and well-aligned with objects.
iConFormer utilizes Input-Conditioned Network (iCoN) to dynamically generate convolutional filters tailored to the input features. These convolutional kernels enable iConFormer to capture detailed spatial features while preserving overall contextual awareness. By focusing on these salient details, iConFormer demonstrates significant improvements in processing complex input data, leading to enhanced accuracy and robustness in dense prediction tasks.
Experiments
Experimental Settings
Datasets and Downstream Tasks.
To evaluate the effectiveness of iConFormer, we conduct comprehensive experiments on both image classification and dense prediction tasks, including monocular depth estimation, semantic segmentation, and instance segmentation. Implementation details can be found in the supplementary material. The datasets used in the experiments are as follows:
  • Image Classification: CIFAR-100 dataset (Krizhevsky, Hinton et al. 2009) consists of 50,000 training images and 10,000 validation images, each with a resolution of 32×32, categorized into 100 classes. The SVHN dataset (Netzer et al. 2011) includes over 600,000 labeled images for digit classification, comprising 73,257 training samples, 26,032 test samples, and 531,131 additional training images. The Food-101 dataset (Bossard, Guillaumin, and Van Gool 2014) contains 101,000 images across 101 food categories, with each category having 750 training samples and 250 test samples.
  • Monocular Depth Estimation: NYU-v2 (Silberman et al. 2012) with diverse indoor scenes and KITTI (Geiger, Lenz, and Urtasun 2012) with high-resolution outdoor driving scenes are benchmark datasets for depth estimation. For experiments, we used the standard splits and evaluate using Root Mean Square Error (RMSE). NYU-v2 images were cropped to 352 × 352 pixels, while KITTI images were cropped to 480 × 480 pixels.
  • Semantic Segmentation: ADE20K (Zhou et al. 2017) is a widely used semantic segmentation dataset with 20,000 training and 2,000 validation images. For our experiments, we utilized UperNet (Xiao et al. 2018) as the framework and evaluated performance using the mean Intersection over Union (mIoU) metric.
  • Instance Segmentation: MS COCO (Lin et al. 2014) is a prominent dataset for instance segmentation, with 118,000 training and 5,000 validation images. We used Cascade Mask R-CNN (Cai and Vasconcelos 2018; He et al. 2017) as a task-specific decoder and measured performance with Average Precision for bounding boxes ($\textrm{AP}_{Box}$) and masks ($\textrm{AP}_{Mask}$).
Pretrained Backbones.
For a fair comparison with FFT baseline and current PEFT methods, we used different pre-trained backbones depending on the tasks. In the semantic segmentation and instance segmentation tasks, Swin Transformer backbones (Liu et al. 2021), pre-trained on ImageNet-22k dataset (Deng et al. 2009), were used (Chen et al. 2019). Specifically, we used the Swin-Large backbone for semantic segmentation and the Swin-Base backbone for instance segmentation. For the monocular depth estimation, we utilized the standard Swin-V2-Base backbone (Liu et al. 2022) pre-trained using the MIM (Xie et al. 2022a). For the classification task, we adopted the ViT backbone (Dosovitskiy et al. 2021) pre-trained using MAE (He et al. 2022a).
Table 1: Performance evaluation of image classification. We report the absolute Top-1 accuracy on the CIFAR-100, SVHN, and Food-101 datasets. indicates a learning rate reduced to 0.1×0.1$\times$0.1 × due to unstable training.
Method Params (M) CIFAR-100 SVHN Food-101
Full 86.04 (100%) 85.90 $97.67^{\dagger}$ $90.09^{\dagger}$
Linear 0.07 (0.08%) 69.83 66.91 69.74
VPT 0.08 (0.09%) 82.44 94.02 82.98
AdaptFormer 1.26 (1.46%) 85.90 96.89 87.61
Adapter 2.46 (2.86%) 86.65 97.09 86.89
iConFormer 1.71 (1.98%) 86.94 97.38 87.97
Table 2: Performance evaluation of monocular depth estimation on the NYU-v2 dataset. The results show comparisons of iConFormer with various parameter-efficient fine-tuning approaches. Results with the symbol $\uparrow$ / $\downarrow$ indicate higher/lower is better.
[…]
Full 86.9 (100%) 0.935 0.991 0.998 0.044 0.304 0.109
Fixed 0 (0%) 0.454 0.748 0.896 0.374 1.018 0.382
Partial-l 12.62 (14.52%) 0.492 0.797 0.928 0.307 0.925 0.346
BitFit 0.14 (0.16%) 0.823 0.969 0.992 0.144 0.466 0.169
LN-tuning 0.05 (0.06%) 0.802 0.963 0.990 0.152 0.496 0.180
Adapter 3.11 (3.45%) 0.901 0.987 0.997 0.104 0.361 0.130
AdaptFormer 1.55 (1.76%) 0.845 0.975 0.994 0.134 0.432 0.159
LoRA 3.08 (3.42%) 0.439 0.733 0.885 0.402 1.045 0.395
iConFormer 2.48 (2.68%) 0.914 0.988 0.998 0.098 0.342 0.122
Baseline methods.
For the image classification task, we used the same set of comparison models as in (Chen et al. 2022), and additionally included the Adapter method from (Houlsby et al. 2019). In the monocular depth estimation, we included comparisons with partial tuning such as BiTFiT (Zaken, Ravfogel, and Goldberg 2021), LN-Tuning (Qi et al. 2022). We also evaluated against Partial-l (Yosinski et al. 2014), which fine-tunes only the final block of the backbone and parameters outside the backbone. For adapter-based methods, we compared with recent approaches, including Adapter (Houlsby et al. 2019), AdaptFormer (Chen et al. 2022), LoRA (Hu et al. 2021), and LoRand (Yin et al. 2023). In the semantic segmentation and instance segmentation tasks, we configured the Adapter model under the same settings as described in (Yin et al. 2023) to ensure a fair comparison. Across all tasks, we also included full fine-tuning (FFT) for performance evaluation.
Main Results
Table 3: Performance evaluation of semantic segmentation on the ADE20K dataset. The results show comparisons of iConFormer with various adapter-based approaches.
[…]
Full 198.58 (100%) 51.5 %
Fixed 0 (0%) 32.21 %
Adapter-B 32.04 (16.13%) 46.23 %
Adapter-T 16.04 (8.08%) 43.51 %
AdaptFormer 2.34 (1.18%) 49.85 %
LoRA 4.57 (2.31%) 49.48 %
LoRand 3.59 (1.84%) 50.67 %
iConFormer 3.26 (1.65%) 50.82%
Table 4: Performance evaluation of instance segmentation on the COCO dataset. The results show comparisons of iConFormer with various adapter-based approaches.
Method Params (M) $\textrm{AP}_{Box}\uparrow$ $\textrm{AP}_{Mask}\uparrow$
Full 89.14 (100%) 51.90 % 45.00 %
Fixed 0 (0%) 15.30 % 10.80 %
Adapter-B 14.38 (16.13%) 46.50 % 40.20 %
Adapter-T 7.20 (8.08%) 43.20 % 38.70 %
AdaptFormer 1.60 (1.79%) 48.90 % 42.50 %
LoRA 3.08 (3.43%) 47.50 % 41.50 %
LoRand 2.39 (2.76%) 51.10 % 44.10 %
iConFormer 2.48 (2.78%) 52.70% 45.70%
Image classification.
We evaluate various fine-tuning approaches using backbones pre-trained via self-supervised learning paradigms (Grill et al. 2020; Chen and He 2021; Choi et al. 2023b, a), as detailed in Table 1. The results demonstrate that iConFormer consistently outperforms linear probing, Visual Prompt Tuning (VPT) methods, and recently proposed adapter-based techniques. Specifically, iConFormer achieves performance improvements of 4.5%, 3.36%, and 4.99% over VPT on the image benchmarks CIFAR-100, SVHN, and Food-101, respectively. Furthermore, when compared to recent adapter-based methods such as Adapter and AdaptFormer, iConFormer shows up to 1.04%, 0.49%, and 1.08% higher accuracy, respectively. Notably, iConFormer also surpasses the full fine-tuning approach by more than 1% Top-1 accuracy on the CIFAR-100 dataset. In summary, iConFormer is not only highly efficient compared to other adapter-based methods but also achieves superior performance with up to 2% of the parameters used in full fine-tuning.
Table 5: Ablation study of input-Conditioned kernel size in iCoN. We perform a quantitative comparison of different kernel sizes on the KITTI dataset. Results with the symbol $\uparrow$ / $\downarrow$ indicate the higher/lower is the better.
[…]
$3\times 3$ 2.48 (2.68%) 0.964 0.996 0.999 0.061 2.302 0.089
$5\times 5$ 4.07 (4.48%) 0.964 0.996 0.999 0.060 2.314 0.090
$7\times 7$ 6.47 (6.93%) 0.963 0.996 0.999 0.060 2.320 0.090
Monocular Depth Estimation.
Table 2 present the performance results for the NYU-v2 datasets. As shown in the tables, iConFormer outperforms other parameter efficient fine-tuning methods in all metrics, with the RMSE value being within 0.04 of the full fine-tuning performance. Moreover, iConFormer shows an RMSE improvement of up to 0.2 compared to partial tuning methods, and an enhancement of up to 0.3 RMSE compared to adapter-based methods such as Adapter, AdaptFormer, and LoRA. These results suggest that iConFormer’s capability to generate and apply input-conditioned kernels significantly contributes to its precision in depth estimation tasks. The performance evaluation on the KITTI dataset is presented in the supplemental material.
Semantic Segmentation.
We present the results of the semantic segmentation task on the ADE20K dataset in Table 3. By fine-tuning fewer than 3.3 million backbone parameters, iConFormer achieves 50.82% mIoU on ADE20K, which is approximately 0.7% lower than Full fine-tuning. Moreover, iConFormer requires fewer parameters for tuning compared to previously proposed adapter-based methods while still achieving superior performance. These results suggest that iConFormer effectively utilizes a limited subset of parameters to capture task-specific information and learn detailed features.
Instance Segmentation.
Table 4 presents the instance segmentation results on the COCO dataset. iConFormer demonstrates significant performance gains by training only 2.78% of the total backbone parameters, surpassing both existing adapter-based methods and full fine-tuning. Specifically, it achieves a 0.8% improvement in $\text{AP}_{box}$ and a 0.7% improvement in $\text{AP}_{mask}$ compared to full fine-tuning. These results effectively reveal the advantages of the proposed method and demonstrate its superiority over full fine-tuning in terms of both storage efficiency and performance. Additionally, these findings suggest that iConFormer optimizes resource utilization through its dynamic kernel approach.
Figure 4: Illustration of the sequential and parallel configurations. The sequential design is shown on the left, and the parallel design on the right.
Ablation Studies
We conducted ablation studies to explore various aspects of iConFormer and identify the key factors that contribute to its performance. All ablation experiments were performed using the dense tasks.
input-Conditioned Kernel Size.
In Table 5, we conducted an ablation study on the KITTI dataset by varying the size of the input-conditioned kernel. We experimented with kernel sizes of 3$\times$3, 5$\times$5, and 7$\times$7. The results demonstrate that iConFormer performs well with a 3$\times$3 kernel across most evaluation metrics. Notably, the RMSE slightly improves as the kernel size decreases, with 3$\times$3 kernel achieving the lowest RMSE. This observation indicates that 3$\times$3 kernel captures essential local features effectively while maintaining computational efficiency. Given that the performance across kernel sizes is quite similar, 3$\times$3 kernel was adopted for its efficiency, providing a balanced trade-off between accuracy and computational cost for the input-conditioned kernels of iCoN.
Table 6: Ablation study of the sequential and parallel configurations on dense prediction tasks. Results are presented in the order of NYU-v2, ADE20K, and COCO datasets, from top to bottom.
[…]
Parallel 2.48 (2.68%) 0.357
Sequential 2.48 (2.68%) 0.342
[…]
Parallel 3.26 (1.65%) 50.25 %
Sequential 3.26 (1.65%) 50.82 %
Configuration Params (M) $\textrm{AP}_{Box}\uparrow$ $\textrm{AP}_{Mask}\uparrow$
Parallel 2.48 (2.78%) 51.0 % 44.4 %
Sequential 2.48 (2.78%) 52.7% 45.7%
iConFormer Configuration.
We investigated the iConFormer configuration by comparing its sequential and parallel instances, as illustrated in Figure 4, where the distinction is based on the placement of the Adapter within the Transformer block. As demonstrated in Table 6, the sequential form of iConFormer significantly outperforms the parallel form across all dense tasks. The reason might be: (1) the sequential design processes each layer’s output in a progressive manner, facilitating deeper feature representations and gradual refinement of complex patterns; (2) the parallel design processes outputs simultaneously, which results in limited inter-layer interaction, weakening the information flow and hindering the model’s capacity to capture intricate features. Therefore, we adopted the sequential design as the default configuration for iConFormer, given its demonstrated superior performance.
Conclusion
In this work, we have presented iConFormer that leverages a parameter-efficient input-conditioned adapter to effectively capture task-specific features and local information with a limited number of learnable parameters in fine-tuning the models for various downstream tasks. iConFormer demonstrates performance comparable to full fine-tuning across image classification, monocular depth estimation, semantic segmentation, and instance segmentation tasks, by tuning only 1.6% to 2.8% of the backbone parameters. iConFormer effectively addresses the limitations of conventional adapter methods and provides superior performances in all tasks. Although our current focus is on vision recognition tasks, we plan to extend iConFormer to other domains such as natural language processing and multi-modal tasks in future work. We anticipate that this extension will inspire further research into efficient adaptation methods and contribute to developing robust solutions across a variety of applications.

### v2

@@
 To secure flexible learning ability on input instances in various downstream tasks, we introduce an input-Conditioned Network (iCoN) in the dynamic adapter that enables instance-level feature transformation.
-To be specific, iCoN generates channel-wise convolutional kernels for each feature and transform it using adaptive convolution process to effectively capture task-specific and fine-grained details tailor to downstream tasks.
-Experimental results demonstrate that by tuning just 1.6% to 2.8% of the Transformer backbone parameters, iConFormer achieves performance comparable to FFT in monocular depth estimation and semantic segmentation, while outperforming it in image classification and instance segmentation.
-Also, the proposed method consistently outperforms recent PEFT methods for all the tasks mentioned above.
+To be specific, iCoN generates channel-wise convolutional kernels for each feature and transform it using adaptive convolution process to effectively capture task-specific details tailored to downstream tasks.
+Experimental results demonstrate that by tuning just 1.6% to 2.8% of the Transformer backbone parameters, iConFormer achieves a performance comparable to FFT in monocular depth estimation and semantic segmentation, while outperforming it in image classification and instance segmentation.
+Additionally, the proposed method consistently outperforms recent PEFT methods for all the tasks mentioned above.
 Figure 1: Quantitative comparison with full fine-tuning (FFT) and PEFT approaches.
-The top graph compares depth prediction errors on NYU-v2 (Silberman et al.
-2012), while the bottom graph presents performance for semantic segmentation on ADE20K (Zhou et al.
-2017) and instance segmentation on COCO (Lin et al.
-2014).
+The top graph compares depth prediction errors on NYU-v2 [55], while the bottom graph presents performance for semantic segmentation on ADE20K [68] and instance segmentation on COCO [45].
 iConFormer consistently outperforms recent PEFT methods in all dense tasks and surpasses FFT in instance segmentation.
-Introduction
+1 Introduction
 As deep neural networks (DNNs) grow increasingly complex, transfer learning—fine-tuning pre-trained models with task-specific data for downstream tasks—has become a widely adopted solution across diverse applications, including image classification, semantic segmentation, and object detection, to name a few.
-For instance, the model consisting of the pre-trained encoder (Xie et al.
-2022b; He et al.
-2022a) and task-specific decoder is fine-tuned, achieving remarkable performance gain when compared to the model trained from scratch.
+For instance, the model consisting of the pre-trained encoder [62, 30] and task-specific decoder is fine-tuned, achieving remarkable performance gain when compared to the model trained from scratch.
 However, training large, complex models separately for each task and dataset is highly inefficient.
-Recently, parameter efficient fine-tuning (PEFT) approaches (Zaken, Ravfogel, and Goldberg 2021; Qi et al.
-2022; Hu et al.
-2021) that maximize the efficiency in terms of training parameters have emerged as an alternative to the above-mentioned full fine tuning (FFT) methodologies, achieving competitive performance even with limited computing resources while simplifying the training processes and deployment.
-This remarkable progress in vision tasks is primarily driven by approaches including prompt tuning (Lester, Al-Rfou, and Constant 2021) and adapter (Houlsby et al.
-2019), which have been successfully applied in natural language processing (NLP) tasks.
-Visual prompt tuning (Jia et al.
-2022) is the first study to explore the potential of prompt tuning in visual recognition tasks, laying the foundation for the prompt tuning in the field of computer vision.
-In addition, the adapter-based PEFT methods (Chen et al.
-2022; Yin et al.
-2023) achieve significant training efficiency by applying the adapter used in NLP to the Vision Transformer (ViT) and its variants (Dosovitskiy et al.
-2021; Liu et al.
-2021, 2022).
-While most PEFT-based approaches yield comparable performance to baseline methods using the FFT in the image classification task, they do not yet provide sufficiently satisfactory performance to compete with the FFT in other complex downstream tasks.
-Additionally, the scalability of prompt-based methods (Jia et al.
-2022; Bahng et al.
-2022) is significantly limited, leading to considerable performance degradation as the number of learnable parameters increases, as reported in (Chen et al.
-2022).
-Adapter-based approaches (Chen et al.
-2022; Hu et al.
-2021; Yin et al.
-2023) incorporate lightweight modules to reduce the number of trainable parameters while maintaining stable performance across a range of trainable parameter scales.
+Recently, parameter efficient fine-tuning (PEFT) approaches [67, 51, 34] that maximize the efficiency in terms of training parameters have emerged as an alternative to the above-mentioned full fine tuning (FFT) methodologies, achieving competitive performance even with limited computing resources while simplifying the training processes and deployment.
+This remarkable progress in vision tasks is primarily driven by approaches including prompt tuning [43] and adapter [33], which have been successfully applied to natural language processing (NLP) tasks.
+Visual Prompt Tuning (VPT) [35] is the first study to explore the potential of prompt tuning in visual recognition tasks, laying the foundation for the prompt tuning in the field of computer vision.
+In addition, the adapter-based PEFT methods [10, 65] achieve significant training efficiency by applying the adapter to the Vision Transformer (ViT) and its variants [22, 47, 48].
+While most PEFT-based approaches yield performance comparable to baseline methods using the FFT in the image classification task, they do not yet provide sufficiently satisfactory performance to compete with the FFT in other complex downstream tasks.
+Furthermore, the scalability of prompt-based methods [35, 3] is significantly limited, leading to considerable performance degradation as the number of learnable parameters (i.e., prompts) increases, as reported in [10].
+In contrast, adapter-based approaches [10, 34, 65] incorporate lightweight modules to reduce the number of trainable parameters, maintaining stable performance on a range of trainable parameter scales.
 However, the adapters always apply the same transformation to input features, ignoring individual characteristics of input instances.
-This may not be an issue when fully tuning DNNs, but it could be a limiting factor in improving performance in adapter-based PEFT methods.
-Namely, the inflexibility with respect to input instances exacerbates the transfer capability of adapter-based models with small learnable parts to downstream tasks, limiting their ability to capture unique and task-specific information.
-Additionally, the Vision Transformer (ViT) backbone (Dosovitskiy et al.
-2021) used in adapter-based models tends to focus on global information rather than fine-grained local details within an image.
-While this limitation can be partially addressed by employing the Swin Transformer (Liu et al.
-2021), which utilizes local attention mechanisms, the constraint on the number of learnable parameters in PEFT still restricts the Swin Transformer’s capability to effectively capture local features.
-Consequently, this negatively affects the performance in dense prediction tasks that require fine-grained information.
-To address the aforementioned issues, we propose a novel PEFT approach, input-Conditioned transFormer, termed iConFormer, that leverages a dynamic adapter where parameters are adjusted input instance-level.
-Unlike recent adapter-based approaches (Chen et al.
-2022; Hu et al.
-2021; Yin et al.
-2023), iConFormer introduces an input-Conditioned Network (iCoN) that dynamically generates the parameters for each input feature in the adapter.
-This approach enables for effectively capturing task-specific and local details while keeping the number of learnable parameters small.
-The effectiveness of our method is evidenced by the quantitative analysis in Figure 1, which shows that iConFormer captures task-specific information more effectively than conventional methods.
+This may not be an issue when fully tuning whole networks, but it could be a limiting factor in improving performance in adapter-based PEFT methods.
+Namely, the inflexibility with respect to input instances exacerbates the transfer capability of adapter-based models with only small learnable parts to downstream tasks, limiting their ability to capture unique and task-specific information.
+Furthermore, the ViT [22] used in adapter-based models tends to focus on global information rather than fine local details within an image.
+While this limitation can be partially addressed by employing the Swin Transformer [47], which utilizes local attention mechanisms, the constraint on the number of learnable parameters in adapter-based approaches still restricts the Swin Transformer’s capability to effectively capture local features (Figure 5).
+Consequently, this negatively affects the performance in dense prediction tasks that require local details.
+To address the aforementioned issues, we propose a novel PEFT approach, input-Conditioned transFormer (iConFormer), which leverages a dynamic adapter where parameters are adjusted at the input instance level, unlike existing adapter-based approaches [10, 34, 65].
+We introduce an input-Conditioned Network (iCoN) that dynamically generates the parameters for each input feature in the adapter.
+This approach enables for effectively capturing task-specific and local details for each instance while keeping the number of learnable parameters small.
+The effectiveness of our method is evidenced by the quantitative analysis in Figure 1.
+Our method achieves performance competitive to the FFT in both classification and dense prediction tasks including monocular depth estimation, semantic segmentation, and instance segmentation with only additional 1.6% to 2.8% backbone parameters.
+Our method even surpasses FFT for the image classification in CIFAR100 [38] and the challenging instance segmentation task on COCO [45].
+Additionally, iConFormer also outperforms conventional PEFT methods for monocular depth estimation task on NYU-v2 [55], demonstrating the effective utilization of pre-trained backbone parameters with additional learnable parameters dynamically finetuned for specific tasks.
 We also analyze the capability to capture fine-grained details by visualizing attention maps in Figure 5.
-iConFormer achieves performance competitive to full fine-tuning (FFT) on both classification and dense prediction tasks including monocular depth estimation, semantic segmentation, and instance segmentation with only additional 1.6% to 2.8% backbone parameters.
-Our method even surpasses FFT for the image classification on CIFAR100 (Krizhevsky, Hinton et al.
-2009) and the challenging instance segmentation task on COCO (Lin et al.
-2014).
-Additionally, iConFormer also outperforms conventional PEFT methods for monocular depth estimation task on NYU-v2 (Silberman et al.
-2012), demonstrating the effective utilization of pre-trained backbone parameters with additional learnable parameters dynamically finetuned for specific tasks.
 In summary, our contributions are threefold:
@@
 (a) FFT, where all parameters are updated during training.
-(b) Our PEFT (iConFormer), where adapters sequentially are attached after the MLP layer in the Transformer.
-Inside the adapters, an Input-Conditioned Network (iCoN) generates input-conditioned convolutional kernels in a channel basis.
-By convolving features with these kernels, iCoN adaptively refines the features in accordance with the specific properties of the input, thereby enhancing the model’s capability to effectively process diverse data in the downstream tasks.
-Related Work
-Transformer in Vision.
-Transformers, initially designed for Natural Language Processing (NLP) tasks such as machine translation (Vaswani 2017) and text generation (Devlin 2018), have achieved significant success in these areas.
-This success has led to a shift towards computer vision, starting with the Vision Transformer (ViT) (Dosovitskiy et al.
-2021).
-Subsequently, various Transformer-based models (Liu et al.
-2021; Xie et al.
-2021; Carion et al.
-2020; Lee et al.
-2022, 2023; Ranftl, Bochkovskiy, and Koltun 2021; Hong et al.
-2022) have achieved notable advancements in tasks including image classification (Krizhevsky, Sutskever, and Hinton 2012), semantic segmentation (Long, Shelhamer, and Darrell 2015; Chen et al.
-2017; Cho et al.
-2024; Kim, Choi, and Min 2022), object detection (Girshick et al.
-2014; Ren et al.
-2016; Redmon et al.
-2016), image restoration (Dong et al.
-2015; Liu, Breuel, and Kautz 2017), and depth estimation (Godard et al.
-2019; Choi et al.
-2021).
-Furthermore, transformers have significantly advanced vision recognition through large-scale pretraining (Chen, Xie, and He 2021; Caron et al.
-2021; He et al.
-2022a).
+(b) Our PEFT (iConFormer), where an dynamic adapter is attached sequentially after the MLP layer in the Transformer.
+Inside the dynamic adapter, an Input-Conditioned Network (iCoN) generates input-conditioned convolutional kernels in a channel basis, which is detailed in Figure 4.
+By convolving features with these kernels, iCoN adaptively refines them in accordance with the specific properties of the input, thereby enhancing the model’s capability to effectively process diverse input data in the downstream tasks.
+2 Related Work
+2.1 Transformer in Vision
+Transformers, initially designed for Natural Language Processing (NLP) tasks such as machine translation [56] and text generation [20], have achieved significant success in these areas.
+This success has led to a shift towards computer vision, starting with the Vision Transformer (ViT) [22].
+Subsequently, various Transformer-based models [47, 60, 6, 40, 41, 52, 32] have achieved notable advancements in tasks including image classification [39], semantic segmentation [49, 9, 13, 37], object detection [24, 54, 53], image restoration [21, 46], and depth estimation [25, 14].
+Furthermore, transformers have significantly advanced vision recognition through large-scale pretraining [12, 7, 30].
 However, their larger size compared to previously prevalent CNN backbones presents challenges for fine-tuning on specific tasks.
 In this context, our work explores methods to adapt pre-trained transformers into target tasks in a more effective and efficient way.
-Parameter Efficient Fine Tuning.
-Parameter Efficient Fine-Tuning (PEFT) methods enable the adaptation of large self-supervised pre-trained models (Xie et al.
-2022c; He et al.
-2022b; Yi et al.
-2022; Choi et al.
-2024a, b) to specific tasks without the need to train the entire model.
-In Natural Language Processing (NLP), notable approaches include adapter methods (Houlsby et al.
-2019), which integrate small modules into the model while keeping the pre-trained parameters frozen, with only the added modules being fine-tuned.
-Additionally, other methods involve tuning specific components, such as bias or normalization layers (Zaken, Ravfogel, and Goldberg 2021; Qi et al.
-2022), utilizing learnable prompt tokens (Lester, Al-Rfou, and Constant 2021), or applying low-rank matrix approximations (Hu et al.
-2021) to update parameters more efficiently.
+2.2 Parameter Efficient Fine Tuning
+Parameter Efficient Fine-Tuning (PEFT) methods enable the adaptation of large pre-trained models [63, 29, 64, 17, 18] to specific tasks without the need to train the entire model.
+In NLP, notable approaches include adapter methods [33], which integrate small learnable modules into the model while keeping the pre-trained parameters frozen, with only the added modules being fine-tuned.
+Additionally, other methods involve tuning specific components such as bias or normalization layers [67, 51], utilizing learnable prompt tokens [43], or applying low-rank matrix approximations [34] to efficiently update parameters.
+Recently, a method has been proposed to improve inference time while maintaining the training parameter efficiency by selectively skipping the computation of less important tokens [42].
 In computer vision, PEFT techniques inspired by NLP have shown significant progress.
-For instance, VPT (Jia et al.
-2022) is the first method to apply prompt tuning approaches to visual classification tasks.
-AdaptFormer (Chen et al.
-2022) employs a parallel adapter framework to enhance the effectiveness of parameter efficient fine-tuning for visual classification.
-KAdaptation (He et al.
-2023) optimizes the adapter using Kronecker products.
-In addition, LoRand (Yin et al.
-2023) employs multi-branch low-rank adapters to achieve impressive performance on dense prediction tasks.
-Our work focuses on adapter methods, proposing new approaches that aim to achieve performance comparable to or exceeding that of full fine-tuning across various downstream tasks.
-Preliminary
-Vision Transformer and its Variants
-Vision Transformer (ViT) (Dosovitskiy et al.
-2021), modified from the Transformer (Vaswani 2017) proposed in NLP, integrates image patches and positional encodings to capture spatial information.
-It consists of a patch embedding layer and multiple sequential encoder blocks, as depicted in Fig.
-2(a).
+VPT [35] is the first method to apply prompt tuning approaches to visual recognition tasks.
+AdaptFormer [10] introduces a parallel adapter framework to enhance the effectiveness of parameter efficient fine-tuning for visual recognition.
+KAdaptation [31] optimizes the adapter using Kronecker products, and SPT [27] selectively allocates trainable parameters to important positions under a specified budget.
+In addition, LoRand [65] employs multi-branch low-rank adapters to achieve impressive performance on dense prediction tasks.
+Our approach is also based on the adapter framework but introduces input-conditioned kernels for instance-specific adaptation, allowing more precise and flexible fine-tuning with a limited number of learnable parameters.
+3 Preliminary
+3.1 Vision Transformer and its Variants
+Vision Transformer (ViT) [22], modified from the Transformer [56] proposed in NLP, integrates image patches and positional encodings to capture spatial information.
+It consists of a patch embedding layer and multiple sequential encoder blocks, as depicted in Figure 2 (a).
 Given a batch of images $x\in\mathbb{R}^{B\times H\times W\times 3}$, the patch embedding layer transforms $x$ into sequential patches $x_{p}\in\mathbb{R}^{B\times M\times(P^{2}C)}$, where $H\times W$ is an image resolution, and $B$ is a batch size.
@@
 The patches are linearly embedded into $D$ dimensions to generate the final input $x_{in}\in\mathbb{R}^{B\times M\times D}$.
-In the Transformer encoder block, $x_{in}$ is first normalized using LayerNorm (Ba, Kiros, and Hinton 2016), and then processed by a multi-head self-attention layer (MHSA).
+In the Transformer encoder block, $x_{in}$ is first normalized using LayerNorm (LN) [2], and then processed by a multi-head self-attention layer (MHSA).
 The output is combined with $x_{in}$ via a residual connection:
@@
 In ViT, the self-attention mechanism captures global features by evaluating relationships between all image patches, enabling a comprehensive understanding of complex dependencies.
-Advancing this approach, Swin Transformer (Liu et al.
-2021, 2022) introduces hierarchical attention with shifted windows, which enhances both computational efficiency and feature representation.
-Other variants (Wang et al.
-2021; Xie et al.
-2021; Li et al.
-2023; Xia et al.
-2024) leverage multi-scale feature extraction for specific vision tasks, demonstrating the adaptability of Transformer models to diverse computer vision challenges.
-PEFT Methods
-Parameter efficient fine-tuning (PEFT) methods, such as prompt tuning (Lester, Al-Rfou, and Constant 2021; Jia et al.
-2022), low-rank adaptation (Hu et al.
-2021), and adapters (Houlsby et al.
-2019; Chen et al.
-2022), are designed to reduce the number of trainable parameters needed for fine-tuning large models.
+Advancing this approach, Swin Transformer [47, 48] introduces hierarchical attention with shifted windows, which enhances both computational efficiency and feature representation.
+Other variants [57, 60, 44, 58] leverage multi-scale feature extraction for specific vision tasks, improving the adaptability of Transformer models.
+3.2 PEFT Methods
+Parameter efficient fine-tuning (PEFT) methods, such as prompt tuning [43, 35], low-rank adaptation [34], and adapters [33, 10], are designed to reduce the number of trainable parameters needed for fine-tuning large models.
 Here, we briefly review adapter-based PEFT methods related to our approach, which will be detailed in the following section.
-Adapters introduce small, trainable modules between the layers of the pre-trained model.
-For instance, if the original model’s transformation is defined as $h=W\cdot x$, adapters modify this operation to $h=W\cdot x+A\cdot g(B\cdot x)$, where $B$ represents the down-projection of the input features, $A$ indicates the up-projection back to the original space, and $g(\cdot)$ denotes a non-linearity.
-These approaches allow for effective task adaptation while minimizing the number of parameters that need to be updated.
-Figure 3: Comparison of Attention Maps from AdaptFormer and iConFormer.
-We visualize the attention maps using attention rollout (Abnar and Zuidema 2020).
-The top row represents input images, and the middle and bottom rows present the attention maps generated by AdaptFormer (Chen et al.
-2022) and iConFormer, respectively, with both using Swin Transformer backbone (Liu et al.
-2021).
-iConFormer more accurately delineates object regions and captures fine-grained semantics, compared to the AdaptFormer.
-Proposed Method
-Conventional adapter-based methods (Houlsby et al.
-2019; Chen et al.
-2022) rely on static parameter tunings, which may constrain their ability to capture diverse input features with high precision.
-To address these limitations, we propose iConFormer, a novel framework incorporating the input-Conditioned Network (iCoN), as illustrated in Figure 2 (b).
-Unlike conventional adapters, iConFormer dynamically generates convolutional kernels tailored to the unique characteristics of each input.
-This dynamic generation facilitates more accurate and flexible feature extraction, introducing locality inductive biases into the pre-trained Transformer encoder.
-This approach enables the model to effectively capture both global and local features within a parameter-efficient framework.
-Consequently, iConFormer not only improves the model’s performance but also ensures a more precise adaptation to varying input conditions.
-Input-Conditioned Network (iCoN)
-The input-Conditioned Network (iCoN) is a key component of the iConFormer framework.
-Inspired by the concept of dynamic filter networks (Jia et al.
-2016), iCoN employs a channel-wise mechanism to generate convolutional kernels that are tailored to the unique characteristics of each input.
+The adapter introduces small, trainable modules between the layers of the pre-trained model.
+It can be integrated in sequential or parallel configurations as shown in Figure 3.
+For instance, if the original architecture processes the features as described in (2), conventional adapters modify this transformation to
+$x_{out}=\begin{cases}\gamma\cdot\text{Up}(\sigma(\text{Down}(\tilde{x})))+\acute{x_{in}},&\text{(Sequential)}\\ \tilde{x}+\gamma\cdot\text{Up}(\sigma(\text{Down}(\acute{x_{in}})))+\acute{x_{in}},&\quad\text{(Parallel)}\\ \end{cases}$ (3)
+where $\text{Down}(\cdot)$ represents the down-projection of the input features, $\text{Up}(\cdot)$ indicates the up-projection back to the original space, and $\sigma$ denotes an activation function.
+Here, $\gamma$ is a weighting factor that adjusts the contribution of the adapter output.
+These approaches allow for task adaptation while minimizing the number of learnable parameters that need to be updated during model training.
+Figure 3: Illustration of the sequential and parallel configurations. The sequential design is shown on the left, and the parallel design on the right.
+4 Proposed Method
+4.1 Motivation and Overview
+Conventional adapter-based methods [33, 10] rely on static parameter-tuning, where the learnable modules such as ‘Up’ and ‘Down’ added to the original transformation are always static with respect to the input feature $\tilde{x}$ or $\acute{x_{in}}$, as described in (3).
+Thus, the same transformation is applied to all instances regardless of instance-specific characteristics, limiting the ability to adapt to input feature distributions in the constrained training environment where only a few number of parameters are tuned for various downstream tasks.
+Namely, in the PEFT that allows for updating only a minimal set of parameters, the static transformations become insufficient for handling diverse input variations, necessitating an instance-level mechanism that dynamically adjusts the learnable parameters conditioned on the input features.
+To this end, we incorporate dynamic kernel generation, enabling instance-aware transformations while maintaining parameter efficiency under PEFT constraints.
+Specifically, we introduce iConFormer, a novel framework incorporating the input-Conditioned Network (iCoN), as illustrated in Figure 2 (b).
+Unlike conventional static adapters, iConFormer dynamically generates input-conditioned convolution kernels tailored to the unique characteristics of each instance, enabling more precise and flexible adaptation.
+By introducing dynamic kernel generation, the proposed method enhances feature extraction while incorporating locality inductive biases into the pretrained Transformer encoder, allowing the model to dynamically capture both global and local features.
+Figure 4: Architecture of input-Conditioned Network (iCoN). The down-projected feature map $\hat{x}$ is used to dynamically generate channel-wise convolution kernels through the iCoN.
+4.2 Input-Conditioned Network (iCoN)
+The iCoN is a key component of the iConFormer framework.
+Inspired by the concept of dynamic filter networks [36], iCoN employs a channel-wise mechanism to dynamically generate convolutional kernels that are tailored to the unique characteristics of each input.
 This approach enhances parameter efficiency while maintaining the flexibility needed to capture diverse features effectively.
@@
 First, the feature $\tilde{x}$ is down-projected to $d$ channels and reshaped from $M$ into the spatial dimensions $\hat{H}\times\hat{W}$, which correspond to $H/P\times W/P$.
-This reshaping effectively rearranges the patches into a spatial grid to produce the feature $\hat{x}\in\mathbb{R}^{B\times\hat{H}\times\hat{W}\times d}$.
+This reshaping rearranges the patches into a spatial grid, producing the feature $\hat{x}\in\mathbb{R}^{B\times\hat{H}\times\hat{W}\times d}$.
 Subsequently, the iCoN module generates dynamic convolutional kernels from the reshaped feature $\hat{x}$.
@@
 $W_{dynamic}=\text{iCoN}(\hat{x}),$ (5)
-where $W_{dynamic}\in\mathbb{R}^{B\times D\times K\times K}$ is the dynamically generated kernel and $K$ is the kernel size.
-For our implementation, we set $K$ to 3 and $d$ to 64 ($d\ll D$).
-The dynamically generated convolution kernel $W_{dynamic}$ is then applied to feature $\hat{x}$, enabling a more precise capture of each input’s distinct features.
+where $W_{dynamic}\in\mathbb{R}^{B\times d\times K\times K}$ is the dynamically generated kernel and $K$ is the kernel size. For our implementation, we set $K$ to 3 and $d$ to 64 ($d\ll D$).
+Figure 4 illustrates the process of dynamically generating convolutional kernels in the iCoN module.
+It first applies spatial average pooling to extract global contextual information from $\hat{x}$.
+The pooled features are then passed through a lightweight transformation, which learns a mapping function to parameterize the convolutional kernel weights.
+Finally, the transformed features are reshaped into the channel-wise convolutional kernel $W_{dynamic}$.
+The convolution kernel generated conditioned on the input feature $\hat{x}$ is then applied to $\hat{x}$ through a channel-wise convolution operation.
 Afterward, the feature is reshaped back to $\mathbb{R}^{B\times M\times d}$, followed by applying a non-linear activation function and up-projection to restore the channel dimension to the original size $D$.
-This process produces the final output of adapter as follows:
+This process produces the final output of the adapter as follows:
 $x_{A}=\text{Up}(\sigma(\text{Reshape}(\hat{x}\otimes W_{dynamic}))),$ (6)
@@
 where $\gamma$ is a weight that adjusts the impact of the adapter features. This weight is a learnable scalar, optimized during training.
-By leveraging dynamically generated convolutional kernels, iConFormer refines the capture of input-specific features compared to conventional adapters, while maintaining parameter efficiency.
-This efficiency is achieved by employing a consistent set of parameters within the iCoN to generate convolutional kernels, even though the kernels are dynamically adapted for each specific input.
-Furthermore, the efficiency is enhanced by generating these dynamic kernels on a channel-wise basis, thereby minimizing the number of tunable parameters while preserving effective adaptation.
-Consequently, iConFormer strikes a balance between effective feature extraction and computational efficiency relative to conventional PEFT methods, addressing their limitations and providing a robust solution for parameter-efficient adaptation.
-Visual Analysis of Local Representation
-To evaluate the performance of capturing both local and global information, Fig.
-3 visualizes the attention maps that offer insights on where the model focuses on.
-The attention maps are generated using attention rollout (Abnar and Zuidema 2020) with AdaptFormer (Chen et al.
-2022) and iConFormer, employing Swin Transformer (Liu et al.
-2021) as the backbone.
-Attention rollout computes token attentions by recursively multiplying attention matrices across layers, revealing how attention is distributed across different regions of the image.
-AdaptFormer adopts a standard pipeline of conducting down-projection, non-linear activation, and up-projection using static weight parameters that are not dynamically adjusted conditioned on input features.
-While the attention maps of AdaptFormer exhibit ambiguous and scattered attention distributions, with limitations in capturing fine-grained local features such as object edges, the attention maps of iConFormer are significantly more focused and well-aligned with objects.
-iConFormer utilizes Input-Conditioned Network (iCoN) to dynamically generate convolutional filters tailored to the input features.
-These convolutional kernels enable iConFormer to capture detailed spatial features while preserving overall contextual awareness.
-By focusing on these salient details, iConFormer demonstrates significant improvements in processing complex input data, leading to enhanced accuracy and robustness in dense prediction tasks.
-Experiments
-Experimental Settings
-Datasets and Downstream Tasks.
-To evaluate the effectiveness of iConFormer, we conduct comprehensive experiments on both image classification and dense prediction tasks, including monocular depth estimation, semantic segmentation, and instance segmentation.
-Implementation details can be found in the supplementary material.
-The datasets used in the experiments are as follows:
-  • Image Classification: CIFAR-100 dataset (Krizhevsky, Hinton et al.
-2009) consists of 50,000 training images and 10,000 validation images, each with a resolution of 32×32, categorized into 100 classes.
-The SVHN dataset (Netzer et al.
-2011) includes over 600,000 labeled images for digit classification, comprising 73,257 training samples, 26,032 test samples, and 531,131 additional training images.
-The Food-101 dataset (Bossard, Guillaumin, and Van Gool 2014) contains 101,000 images across 101 food categories, with each category having 750 training samples and 250 test samples.
-  • Monocular Depth Estimation: NYU-v2 (Silberman et al.
-2012) with diverse indoor scenes and KITTI (Geiger, Lenz, and Urtasun 2012) with high-resolution outdoor driving scenes are benchmark datasets for depth estimation.
-For experiments, we used the standard splits and evaluate using Root Mean Square Error (RMSE).
-NYU-v2 images were cropped to 352 × 352 pixels, while KITTI images were cropped to 480 × 480 pixels.
-  • Semantic Segmentation: ADE20K (Zhou et al.
-2017) is a widely used semantic segmentation dataset with 20,000 training and 2,000 validation images.
-For our experiments, we utilized UperNet (Xiao et al.
-2018) as the framework and evaluated performance using the mean Intersection over Union (mIoU) metric.
-  • Instance Segmentation: MS COCO (Lin et al.
-2014) is a prominent dataset for instance segmentation, with 118,000 training and 5,000 validation images.
-We used Cascade Mask R-CNN (Cai and Vasconcelos 2018; He et al.
-2017) as a task-specific decoder and measured performance with Average Precision for bounding boxes ($\textrm{AP}_{Box}$) and masks ($\textrm{AP}_{Mask}$).
-Pretrained Backbones.
-For a fair comparison with FFT baseline and current PEFT methods, we used different pre-trained backbones depending on the tasks.
-In the semantic segmentation and instance segmentation tasks, Swin Transformer backbones (Liu et al.
-2021), pre-trained on ImageNet-22k dataset (Deng et al.
-2009), were used (Chen et al.
-2019).
-Specifically, we used the Swin-Large backbone for semantic segmentation and the Swin-Base backbone for instance segmentation.
-For the monocular depth estimation, we utilized the standard Swin-V2-Base backbone (Liu et al.
-2022) pre-trained using the MIM (Xie et al.
-2022a).
-For the classification task, we adopted the ViT backbone (Dosovitskiy et al.
-2021) pre-trained using MAE (He et al.
-2022a).
+By dynamically generating convolutional kernels, iCoN allows the model to effectively adapt to the input structure while enriching the feature representation with both global context and local details.
+Here, it should be noted that the standard convolution operation can also be applied to the Transformer encoder to inject the locality, but the dynamic adaptation to the input instance is more crucial to the PEFT framework for various downstream tasks.
+In contrast to standard convolutional kernels, iCoN dynamically adjusts its kernels based on the spatial structure of each input, resulting in more precise and context-aware feature extraction.
+This mechanism allows iCoN to capture high-frequency variations and subtle patterns that the standard convolution layers struggle to represent effectively.
+This is also validated in experiments (Table 7).
+It is worthy of noting that while our approach generates dynamic kernels, it remains computationally comparable to conventional adapter-based PEFT methods [10, 33].
+This efficiency is primarily attributed to two factors.
+First, while iCoN generates dynamic kernels, the network responsible for the kernel generation retains static learnable parameters, ensuring that only the output kernels adapt to instance without significantly increasing parameter overhead.
+Second, to further reduce computational complexity while preserving adaptive capacity, iCoN employs the channel-wise convolutions instead of the original convolutions, as shown in Figure 4.
+Figure 5: Comparison of Attention Maps from AdaptFormer and iConFormer.
+We visualize the attention maps using attention rollout [1].
+The top row represents input images, and the middle and bottom rows present the attention maps generated by AdaptFormer [10] and iConFormer, respectively, with both using Swin Transformer backbone [47].
+iConFormer more accurately delineates object regions and captures fine-grained semantics, compared to the AdaptFormer.
+4.3 Visual Analysis of Local Representation
+To evaluate the effectiveness of capturing both local and global information, Figure 5 visualizes the attention maps that provide insight where the model focuses.
+The attention maps are generated using attention rollout [1] with AdaptFormer [10] and iConFormer, employing the Swin Transformer [47] as the backbone.
+Attention rollout computes token attentions by recursively multiplying attention matrices across layers, revealing how attention is distributed across different regions of an input image.
+AdaptFormer adopts a standard pipeline consisting of down-projection, non-linear activation, and up-projection based on static weight parameters that are not dynamically adjusted conditioned on input features.
+While the attention maps of AdaptFormer exhibit ambiguous and scattered attention distributions with limitations in precisely capturing local features such as object edges, the attention maps of iConFormer are significantly more focused and better aligned with objects.
+The iCoN used in the iConFormer dynamically generates convolutional kernels tailored to the input features, enabling the iConFormer to capture detailed spatial features while preserving overall contextual awareness.
+By focusing on these salient details, the proposed method demonstrates significant improvements in processing complex input data, leading to enhanced accuracy and robustness in dense prediction tasks.
+5 Experiments
 Table 1: Performance evaluation of image classification.
@@
 indicates a learning rate reduced to 0.1×0.1$\times$0.1 × due to unstable training.
-Method Params (M) CIFAR-100 SVHN Food-101
-Full 86.04 (100%) 85.90 $97.67^{\dagger}$ $90.09^{\dagger}$
-Linear 0.07 (0.08%) 69.83 66.91 69.74
-VPT 0.08 (0.09%) 82.44 94.02 82.98
-AdaptFormer 1.26 (1.46%) 85.90 96.89 87.61
-Adapter 2.46 (2.86%) 86.65 97.09 86.89
-iConFormer 1.71 (1.98%) 86.94 97.38 87.97
+Additionally, FLOPs are measured with a batch size of 2.
+Method Params (M) FLOPs (G) CIFAR-100 SVHN Food-101
+Full 86.04 (100%) 35.16 85.90 97.67 90.09
+Linear 0.07 (0.08%) 35.16 69.83 66.91 69.74
+VPT [35] 0.08 (0.09%) 35.35 82.44 94.02 82.98
+AdaptFormer [10] 1.26 (1.46%) 35.63 85.90 96.89 87.61
+Adapter [33] 2.46 (2.86%) 36.13 86.65 97.09 86.89
+iConFormer 1.71 (1.98%) 35.65 86.94 97.38 87.97
+5.1 Experimental Settings
 Table 2: Performance evaluation of monocular depth estimation on the NYU-v2 dataset.
@@
 Fixed 0 (0%) 0.454 0.748 0.896 0.374 1.018 0.382
-Partial-l 12.62 (14.52%) 0.492 0.797 0.928 0.307 0.925 0.346
-BitFit 0.14 (0.16%) 0.823 0.969 0.992 0.144 0.466 0.169
-LN-tuning 0.05 (0.06%) 0.802 0.963 0.990 0.152 0.496 0.180
-Adapter 3.11 (3.45%) 0.901 0.987 0.997 0.104 0.361 0.130
-AdaptFormer 1.55 (1.76%) 0.845 0.975 0.994 0.134 0.432 0.159
-LoRA 3.08 (3.42%) 0.439 0.733 0.885 0.402 1.045 0.395
+Partial-l [66] 12.62 (14.52%) 0.492 0.797 0.928 0.307 0.925 0.346
+BitFit [67] 0.14 (0.16%) 0.823 0.969 0.992 0.144 0.466 0.169
+LN-tuning [51] 0.05 (0.06%) 0.802 0.963 0.990 0.152 0.496 0.180
+Adapter [33] 3.11 (3.45%) 0.901 0.987 0.997 0.104 0.361 0.130
+AdaptFormer [10] 1.55 (1.76%) 0.845 0.975 0.994 0.134 0.432 0.159
+LoRA [34] 3.08 (3.42%) 0.439 0.733 0.885 0.402 1.045 0.395
 iConFormer 2.48 (2.68%) 0.914 0.988 0.998 0.098 0.342 0.122
-Baseline methods.
-For the image classification task, we used the same set of comparison models as in (Chen et al.
-2022), and additionally included the Adapter method from (Houlsby et al.
-2019).
-In the monocular depth estimation, we included comparisons with partial tuning such as BiTFiT (Zaken, Ravfogel, and Goldberg 2021), LN-Tuning (Qi et al.
-2022).
-We also evaluated against Partial-l (Yosinski et al.
-2014), which fine-tunes only the final block of the backbone and parameters outside the backbone.
-For adapter-based methods, we compared with recent approaches, including Adapter (Houlsby et al.
-2019), AdaptFormer (Chen et al.
-2022), LoRA (Hu et al.
-2021), and LoRand (Yin et al.
-2023).
-In the semantic segmentation and instance segmentation tasks, we configured the Adapter model under the same settings as described in (Yin et al.
-2023) to ensure a fair comparison.
-Across all tasks, we also included full fine-tuning (FFT) for performance evaluation.
-Main Results
 Table 3: Performance evaluation of semantic segmentation on the ADE20K dataset. The results show comparisons of iConFormer with various adapter-based approaches.
 […]
-Full 198.58 (100%) 51.5 %
+Full 198.58 (100%) 51.50 %
 Fixed 0 (0%) 32.21 %
-Adapter-B 32.04 (16.13%) 46.23 %
-Adapter-T 16.04 (8.08%) 43.51 %
-AdaptFormer 2.34 (1.18%) 49.85 %
-LoRA 4.57 (2.31%) 49.48 %
-LoRand 3.59 (1.84%) 50.67 %
-iConFormer 3.26 (1.65%) 50.82%
+Adapter-B [33] 32.04 (16.13%) 46.23 %
+Adapter-T [33] 16.04 (8.08%) 43.51 %
+AdaptFormer [10] 2.34 (1.18%) 49.85 %
+LoRA [34] 4.57 (2.31%) 49.48 %
+LoRand [65] 3.59 (1.84%) 50.67 %
+iConFormer 3.26 (1.65%) 51.17%
 Table 4: Performance evaluation of instance segmentation on the COCO dataset. The results show comparisons of iConFormer with various adapter-based approaches.
@@
 Fixed 0 (0%) 15.30 % 10.80 %
-Adapter-B 14.38 (16.13%) 46.50 % 40.20 %
-Adapter-T 7.20 (8.08%) 43.20 % 38.70 %
-AdaptFormer 1.60 (1.79%) 48.90 % 42.50 %
-LoRA 3.08 (3.43%) 47.50 % 41.50 %
-LoRand 2.39 (2.76%) 51.10 % 44.10 %
-iConFormer 2.48 (2.78%) 52.70% 45.70%
-Image classification.
-We evaluate various fine-tuning approaches using backbones pre-trained via self-supervised learning paradigms (Grill et al.
-2020; Chen and He 2021; Choi et al.
-2023b, a), as detailed in Table 1.
-The results demonstrate that iConFormer consistently outperforms linear probing, Visual Prompt Tuning (VPT) methods, and recently proposed adapter-based techniques.
-Specifically, iConFormer achieves performance improvements of 4.5%, 3.36%, and 4.99% over VPT on the image benchmarks CIFAR-100, SVHN, and Food-101, respectively.
-Furthermore, when compared to recent adapter-based methods such as Adapter and AdaptFormer, iConFormer shows up to 1.04%, 0.49%, and 1.08% higher accuracy, respectively.
-Notably, iConFormer also surpasses the full fine-tuning approach by more than 1% Top-1 accuracy on the CIFAR-100 dataset.
-In summary, iConFormer is not only highly efficient compared to other adapter-based methods but also achieves superior performance with up to 2% of the parameters used in full fine-tuning.
+Adapter-B [33] 14.38 (16.13%) 46.50 % 40.20 %
+Adapter-T [33] 7.20 (8.08%) 43.20 % 38.70 %
+AdaptFormer [10] 1.60 (1.79%) 48.90 % 42.50 %
+LoRA [34] 3.08 (3.43%) 47.50 % 41.50 %
+LoRand [65] 2.39 (2.76%) 51.10 % 44.10 %
+iConFormer 2.48 (2.78%) 52.90% 45.90%
+Datasets and Downstream Tasks
+To evaluate the performance of iConFormer, we conducted comprehensive experiments on both image classification and dense prediction tasks, including monocular depth estimation, semantic segmentation, and instance segmentation.
+Implementation details can be found in the supplementary material.
+The datasets used in the experiments are as follows:
+  • Image Classification: CIFAR-100 dataset [38] consists of 50,000 training images and 10,000 validation images, each with a resolution of 32×32, categorized into 100 classes.
+The SVHN dataset [50] includes over 600,000 labeled images for digit classification, comprising 73,257 training samples, 26,032 test samples, and 531,131 additional training images.
+The Food-101 dataset [4] contains 101,000 images across 101 food categories, with each category having 750 training samples and 250 test samples.
+  • Monocular Depth Estimation: NYU-v2 [55] with diverse indoor scenes and KITTI [23] with high-resolution outdoor driving scenes are benchmark datasets for depth estimation.
+For experiments, we used the standard splits and evaluate using Root Mean Square Error (RMSE).
+NYU-v2 images were cropped to 352 × 352 pixels, while KITTI images were cropped to 480 × 480 pixels.
+  • Semantic Segmentation: ADE20K [68] is a widely used semantic segmentation dataset with 20,000 training and 2,000 validation images.
+For our experiments, we utilized UperNet [59] as the framework and evaluated performance using the mean Intersection over Union (mIoU) metric.
+  • Instance Segmentation: MS COCO [45] is a prominent dataset for instance segmentation, with 118,000 training and 5,000 validation images.
+We used Cascade Mask R-CNN [5, 28] as a task-specific decoder and measured performance with Average Precision for bounding boxes ($\textrm{AP}_{Box}$) and masks ($\textrm{AP}_{Mask}$).
+Pretrained Backbones
+For a fair comparison with FFT baseline and current PEFT methods, we used different pre-trained backbones depending on the tasks.
+In the semantic segmentation and instance segmentation tasks, Swin Transformer backbones [47], pre-trained on ImageNet-22k dataset [19], were used [8].
+Specifically, we used the Swin-Large backbone for semantic segmentation and the Swin-Base backbone for instance segmentation.
+For the monocular depth estimation, we utilized the standard Swin-V2-Base backbone [48] pre-trained using the MIM [61].
+For the classification task, we adopted the ViT backbone [22] pre-trained using MAE [30].
+Baseline Methods
+For the image classification task, we used the same set of comparison models as [10], and additionally included the Adapter method from [33].
+In the monocular depth estimation, we included comparisons with partial tuning methods such as BiTFiT [67], LN-Tuning [51].
+We also evaluated against Partial-l [66], which fine-tunes only the final block of the backbone and parameters outside the backbone.
+For comparison with adapter-based methods, we included recent approaches such as Adapter [33], AdaptFormer [10], LoRA [34], and LoRand [65].
+In the semantic segmentation and instance segmentation tasks, we configured the Adapter [33] following [65] for a fair comparison, setting the intermediate layer dimension to half of the input dimension for ‘Adapter-B’ and to a quarter for ‘Adapter-T’.
+Additionally, we included ‘Fixed’ in all dense prediction tasks, freezing the pre-trained Transformer encoder while training other parts of the architecture (i.e., task decoder).
+Across all tasks, we also included ‘Full’, which indicates full fine-tuning (FFT) as an upper bound on performance.
+5.2 Main Results
+Image Classification
+We evaluated various fine-tuning approaches using ViT backbone [22] pre-trained via self-supervised learning paradigms [26, 11, 16, 15], as detailed in Table 1.
+The results demonstrate that the iConFormer consistently outperforms linear probing, Visual Prompt Tuning (VPT) methods, and recently proposed adapter-based techniques.
+Specifically, the iConFormer achieves performance improvements of 4.5%, 3.36%, and 4.99% over VPT on the image benchmarks CIFAR-100, SVHN, and Food-101, respectively.
+Furthermore, when compared to recent adapter-based methods such as Adapter [33] and AdaptFormer [10], iConFormer shows up to 1.04%, 0.49%, and 1.08% higher accuracy, respectively.
+Notably, iConFormer also surpasses the FFT approach by more than 1% Top-1 accuracy on the CIFAR-100 dataset.
+Additionally, while consistently delivering better performance, iConFormer demonstrates computational efficiency with 35.65 GFLOPs, which is comparable to AdaptFormer (35.63G) and lower than Adapter (36.13G).
+In summary, our approach outperforms recent adapter-based approaches with similar computational efficiency, and provides comparable performance to the FFT despite tuning only 2% of the parameters used in the FFT.
+Monocular Depth Estimation
+Table 2 presents the performance results for the NYU-v2 datasets.
+As shown in the tables, the iConFormer outperforms other PEFT methods in all metrics, with the RMSE value being within 0.04 of the FFT performance.
+Moreover, the iConFormer shows an RMSE improvement of up to 0.2 compared to partial tuning methods, and an enhancement of up to 0.3 RMSE compared to adapter-based methods such as Adapter [33], AdaptFormer [10], and LoRA [34].
+These results suggest that iConFormer’s capability to generate and apply input-conditioned kernels significantly contributes to the performance in the monocular depth estimation task.
+Additional results on the KITTI dataset are presented in the supplemental material.
+Semantic Segmentation
+We present the results of the semantic segmentation task on the ADE20K dataset [68] in Table 3.
+By fine-tuning fewer than 3.3 million backbone parameters, the proposed method achieves 51.17% mIoU on ADE20K, which is about 0.3% lower than the FFT.
+Moreover, the iConFormer requires fewer learnable parameters compared to most adapter-based methods while still achieving superior performance.
+These results suggest that iConFormer effectively utilizes a limited subset of parameters to capture task-specific information and learn detailed features.
+Instance Segmentation
+Table 4 presents the instance segmentation results on the COCO dataset.
+Our method demonstrates significant performance gains by training only 2.78% of the total backbone parameters, surpassing both existing adapter-based methods and FFT.
+Specifically, it achieves 1.0% improvement in $\text{AP}_{box}$ and 0.9% improvement in $\text{AP}_{mask}$ compared to the FFT.
+These results reveal the advantages of the proposed method and demonstrate its superiority over the FFT in terms of both storage efficiency and performance.
+Additionally, these findings suggest that iConFormer optimizes resource utilization through its dynamic kernel approach.
+Table 5: Ablation study of the sequential and parallel configurations on dense prediction tasks. Results are presented in the order of NYU-v2, ADE20K, and COCO, from left to right.
+Configuration RMSE ↓ mIoU ↑ $\textrm{AP}_{Box}\uparrow$ $\textrm{AP}_{Mask}\uparrow$
+Parallel 0.357 50.85 % 51.20 % 44.60 %
+Sequential 0.342 51.17 % 52.90 % 45.90 %
 Table 6: Ablation study of input-Conditioned kernel size in iCoN.
-We perform a quantitative comparison of different kernel sizes on the KITTI dataset.
+We perform a quantitative comparison of different kernel sizes across dense prediction datasets.
 Results with the symbol $\uparrow$ / $\downarrow$ indicate the higher/lower is the better.
+Kernel Size KITTI ADE20K COCO
+Params (M) RMSE ↓ Params (M) mIoU ↑ Params (M) $\textrm{AP}_{Box}\uparrow$ $\textrm{AP}_{Mask}\uparrow$
+3 $\times$ 3 2.48 (2.68%) 2.302 3.26 (1.65%) 51.17 % 2.48 (2.78%) 52.90 % 45.90 %
+5 $\times$ 5 4.07 (4.48%) 2.314 4.86 (2.43%) 50.95 % 4.07 (4.49%) 52.60 % 45.70 %
+7 $\times$ 7 6.47 (6.93%) 2.320 7.26 (3.59%) 50.87 % 6.47 (6.94%) 52.60 % 45.70 %
+Table 7: Ablation study on the effect of the Input-Conditioned Kernel in dense prediction tasks.
+Results are presented in the order of NYU-v2, ADE20K, and COCO datasets, from top to bottom.
+Both kernel types uses a 3$\times$3 kernel size for all experiments.
 […]
-$3\times 3$ 2.48 (2.68%) 0.964 0.996 0.999 0.061 2.302 0.089
-$5\times 5$ 4.07 (4.48%) 0.964 0.996 0.999 0.060 2.314 0.090
-$7\times 7$ 6.47 (6.93%) 0.963 0.996 0.999 0.060 2.320 0.090
-Monocular Depth Estimation.
-Table 2 present the performance results for the NYU-v2 datasets.
-As shown in the tables, iConFormer outperforms other parameter efficient fine-tuning methods in all metrics, with the RMSE value being within 0.04 of the full fine-tuning performance.
-Moreover, iConFormer shows an RMSE improvement of up to 0.2 compared to partial tuning methods, and an enhancement of up to 0.3 RMSE compared to adapter-based methods such as Adapter, AdaptFormer, and LoRA.
-These results suggest that iConFormer’s capability to generate and apply input-conditioned kernels significantly contributes to its precision in depth estimation tasks.
-The performance evaluation on the KITTI dataset is presented in the supplemental material.
-Semantic Segmentation.
-We present the results of the semantic segmentation task on the ADE20K dataset in Table 3.
-By fine-tuning fewer than 3.3 million backbone parameters, iConFormer achieves 50.82% mIoU on ADE20K, which is approximately 0.7% lower than Full fine-tuning.
-Moreover, iConFormer requires fewer parameters for tuning compared to previously proposed adapter-based methods while still achieving superior performance.
-These results suggest that iConFormer effectively utilizes a limited subset of parameters to capture task-specific information and learn detailed features.
-Instance Segmentation.
-Table 4 presents the instance segmentation results on the COCO dataset.
-iConFormer demonstrates significant performance gains by training only 2.78% of the total backbone parameters, surpassing both existing adapter-based methods and full fine-tuning.
-Specifically, it achieves a 0.8% improvement in $\text{AP}_{box}$ and a 0.7% improvement in $\text{AP}_{mask}$ compared to full fine-tuning.
-These results effectively reveal the advantages of the proposed method and demonstrate its superiority over full fine-tuning in terms of both storage efficiency and performance.
-Additionally, these findings suggest that iConFormer optimizes resource utilization through its dynamic kernel approach.
-Figure 4: Illustration of the sequential and parallel configurations. The sequential design is shown on the left, and the parallel design on the right.
-Ablation Studies
-We conducted ablation studies to explore various aspects of iConFormer and identify the key factors that contribute to its performance. All ablation experiments were performed using the dense tasks.
-input-Conditioned Kernel Size.
-In Table 5, we conducted an ablation study on the KITTI dataset by varying the size of the input-conditioned kernel.
-We experimented with kernel sizes of 3$\times$3, 5$\times$5, and 7$\times$7.
-The results demonstrate that iConFormer performs well with a 3$\times$3 kernel across most evaluation metrics.
-Notably, the RMSE slightly improves as the kernel size decreases, with 3$\times$3 kernel achieving the lowest RMSE.
-This observation indicates that 3$\times$3 kernel captures essential local features effectively while maintaining computational efficiency.
-Given that the performance across kernel sizes is quite similar, 3$\times$3 kernel was adopted for its efficiency, providing a balanced trade-off between accuracy and computational cost for the input-conditioned kernels of iCoN.
-Table 6: Ablation study of the sequential and parallel configurations on dense prediction tasks. Results are presented in the order of NYU-v2, ADE20K, and COCO datasets, from top to bottom.
+     Standard Conv      2.44 (2.64%)      1.029
+     input-Conditioned Conv      2.48 (2.68%)      0.342
 […]
-Parallel 2.48 (2.68%) 0.357
-Sequential 2.48 (2.68%) 0.342
-[…]
-Parallel 3.26 (1.65%) 50.25 %
-Sequential 3.26 (1.65%) 50.82 %
-Configuration Params (M) $\textrm{AP}_{Box}\uparrow$ $\textrm{AP}_{Mask}\uparrow$
-Parallel 2.48 (2.78%) 51.0 % 44.4 %
-Sequential 2.48 (2.78%) 52.7% 45.7%
-iConFormer Configuration.
-We investigated the iConFormer configuration by comparing its sequential and parallel instances, as illustrated in Figure 4, where the distinction is based on the placement of the Adapter within the Transformer block.
-As demonstrated in Table 6, the sequential form of iConFormer significantly outperforms the parallel form across all dense tasks.
+     Standard Conv      3.25 (1.64%)      50.02 %
+     input-Conditioned Conv      3.26 (1.65%)      51.17 %
+Kernel Type Params (M) $\textrm{AP}_{Box}\uparrow$ $\textrm{AP}_{Mask}\uparrow$
+Standard Conv 2.46 (2.76%) 50.20 % 43.50 %
+input-Conditioned Conv 2.48 (2.78%) 52.90 % 45.90 %
+5.3 Ablation Studies
+We conducted ablation studies to explore various aspects of the iConFormer and identify key factors that contribute to its performance.
+All ablation experiments were conducted using the dense predictions tasks.
+iConFormer Configuration
+We investigated the performance by comparing sequential and parallel configurations, as illustrated in Figure 3, where the distinction is based on the placement of the Adapter within the Transformer block.
+As demonstrated in Table 5, the sequential form significantly outperforms the parallel form for all dense tasks.
 The reason might be: (1) the sequential design processes each layer’s output in a progressive manner, facilitating deeper feature representations and gradual refinement of complex patterns; (2) the parallel design processes outputs simultaneously, which results in limited inter-layer interaction, weakening the information flow and hindering the model’s capacity to capture intricate features.
 Therefore, we adopted the sequential design as the default configuration for iConFormer, given its demonstrated superior performance.
-Conclusion
+input-Conditioned Kernel Size
+In Table 6, we present an ablation study on the size of the input-conditioned convolution on dense prediction tasks.
+Experiments with kernel sizes of 3$\times$3, 5$\times$5, and 7$\times$7 demonstrate that the 3×3 kernel consistently achieves competitive results with a relatively small number of parameters.
+Notably, on the KITTI dataset, the RMSE slightly improves as the kernel size decreases, with 3$\times$3 kernel achieving the lowest RMSE.
+Similarly, on the ADE20K and COCO datasets, the 3$\times$3 kernel consistently outperforms the 5$\times$5 and 7$\times$7 variants in both mIoU and AP.
+These results indicate that 3$\times$3 kernel captures essential local features effectively while maintaining computational efficiency.
+Given that the performance across kernel sizes is quite similar, 3$\times$3 kernel was adopted for its efficiency, providing a balanced trade-off between accuracy and computational cost for the input-conditioned kernels of iCoN.
+Effect of input-Conditioned Kernel
+We investigated the effect of using the input-conditioned convolution $W_{dynamic}$ of (6) in the iCoN for dense prediction tasks.
+In Table 7, we compared the performance when using the standard convolution and the input-conditioned convolution for the NYU-v2 dataset (monocular depth estimation), the ADE20K dataset (semantic segmentation), and the COCO dataset (instance segmentation).
+For a fair comparison, we set to $3\times 3$ kernel for both cases, ensuring the same local receptive field.
+The input-conditioned kernel consistently outperforms the standard variant for all tasks, improving mIoU by about 1.2% on ADE20K and AP by about 2.5% on COCO, and reducing RMSE by about 0.7 on NYU-v2.
+To further analyze the performance of the standard convolution, we extended the comparison to existing PEFT methods.
+In Table 2, most existing PEFT methods achieve a lower RMSE than the standard convolution.
+In Table 3 and 4, compared to LoRand [65], the standard convolution achieves approximately 0.6% lower mIoU on ADE20K and about 1% lower AP on COCO, respectively.
+This indicates that simply applying the standard convolution to our framework is not so effective and the adaptive nature of the input-conditioned convolution, where kernel weights are dynamically modulated for input features, is more crucial to capture local details in dense prediction tasks.
+6 Conclusion
 In this work, we have presented iConFormer that leverages a parameter-efficient input-conditioned adapter to effectively capture task-specific features and local information with a limited number of learnable parameters in fine-tuning the models for various downstream tasks.